OpenStudy (anonymous):

n+5(n-1)=7 Do I distribute 5 or 5n first?

OpenStudy (anonymous):

5

jimthompson5910 (jim_thompson5910):

You distribute the 5 first

OpenStudy (anonymous):

Thanks SO MUCH I was so confused!

OpenStudy (anonymous):

n+5(n-1)=7 n+5n-5=7 n+5n = 7 +5 n+5n = 12 6n = 12 6n/6 = 12/6 n = 2

OpenStudy (anonymous):

Thanks!!!